A coupon bond that pays interest of $55 annually has a par value of $1,000, matures in 5 years, and is selling today at a $74.00 discount from par value. The current yield on this bond is _________.
Answer #1

CURRENT YIELD = COUPON INTEREST/CURRENT PRICE

COUPON INTEREST = 55, CURRENT PRICE = 1000 - 74 = 926

CURRENT YIELD = 55/926 = 0.059395 = 5.94%

ANSWER : 5.94% (THUMBS UP PLEASE)